# Limiting noise and outliers in the MICROSCOPE space test of the weak equivalence principle

### v3

@@ II Main features of the MICROSCOPE accelerometers to test WEP/UFF in space
A higher stiffness results in a higher level of random thermal noise from internal damping (Sec. III). It also affects the control of the TM which keeps it motionless at a selected capacitive zero so that the voltage applied yields (through a scale factor in $\rm ms^{-2}/V$ established by calibration) the acceleration which caused the motion.
In absence of control the TM will move, at any time, to an equilibrium position between the tidal force from Earth and the restoring force of the wire, at an elongation $X_{eq}$ from the zero of the elastic force, $0_{k}$. The control acts to nullify this displacement in order to keep the TM at the selected capacitive zero, $0_{C}$. If the two zeros do not coincide, the TM in $0_{C}$ is subjected to a constant elastic force that the control must nullify by applying a constant voltage, in addition to the voltage required to nullify the tidal acceleration (at $2\nu_{{}_{EP}}$) [^2] With no wire and negligible electrostatic stiffness, the zero of each TM control (dictated by geometry, construction, mounting and electronics) does not coincide with the zero of the largest force on the test mass, due to the Earth’s tide. Applying a constant bias is therefore inevitable. Hence, any fluctuation of the bias translates into a fluctuation of the zero of the acceleration measurement and –if it occurs in the frequency region of the WEP violation signal– it contributes to the limiting noise. There is no such issue if the experiment is designed as an intrinsically differential, null experiment (no violation, no signal) as in torsion balances Adel2008 ; AdelCQGfocusissue2012 and the proposed GG space experiment PerBraginskyPLA2018 .
@@
The acceleration biases and their time evolution are reported, for all four accelerometers in ChhunCQG2022-5  Fig. 9. For SUEP the bias of the outer TM is $\simeq 10^{-6}\,\rm ms^{-2}$ and almost one order of magnitude smaller for the inner TM. With different offsets between their respective zeros, different masses and somewhat different stiffness, the acceleration bias is different on the two test masses of each sensor unit and the larger one dominates the acceleration difference, which is the physical quantity of interest. The time series for SUEP session #218 reported by RodriguesRome2018 , Slide 13, confirms that they are at the level of $10^{-6}\,\rm ms^{-2}$. For SUREF, the biases are of a few $10^{-7}\,\rm ms^{-2}$hhunCQG2022-5 . For such biases orders of magnitude above the target not to impair the measurement, it is mandatory that they remain constant. This will not be the case in the presence of patch charges on the surface of the TMs and their electrodes, as they are known to give rise to low frequency fluctuating potentials (Sec. III).
III Limiting noise, patch charges and fluctuating zero
@@
Ever since the preliminary results MS-PRL2017 , the residual random acceleration differences in the frequency region of the signal were attributed to thermal noise from internal damping. Reaching the level of thermal noise is indeed a remarkable achievement.
@@
In this case the line which best fits the spectral density of the limiting noise decreases with the frequency as $1/\sqrt{\nu}$ and reads SaulsonInternalDamping :
@@
For a measurement session of duration $t_{int}$, as long as the acceleration noise is random, and therefore is bound to decrease as the square root of the integration time, the value of (2) at $\nu_{{}_{EP}}$ sets the limit $\delta$ to the violation signal established by the measurement session, according to the relationship:
@@
where $g_{{}_{drive}}=7.9\,\rm ms^{-2}$ is the gravitational acceleration from Earth at the spacecraft orbital distance, $\delta\cdot g_{drive}$ is a tiny acceleration difference between the TMs at the known signal frequency $\nu_{{}_{EP}}$, and $\delta$ is a solve-for parameter resulting from the confrontation between the predicted acceleration differences and their measured values (as defined in TouboulCQG2022-9 , Eq.(6)). [^3] Over a synodic period relative to the Earth $\delta$ changes sign because the phase of WEP violation goes from zero to $\pi$ (depending on whether the differential acceleration vector is pointing towards the CoM of the Earth or away from it; the sign in either case depends on which mass would be more attracted from Earth than the other, which is not known). Note that $\delta$ is never zero because thermal noise goes to zero only for an infinite measurement time. Should a WEP violation signal be buried by fluctuations around the best fitting line (2), the smaller their size, the shorter the additional integration time that would be required for it to emerge and be detected. These fluctuations yield the value $\sigma$ of the resulting $\delta\pm\sigma$ for the session.
@@
Early on, the spin rate was designed to be between $3$ and $5$ times the orbital frequency $\nu_{orb}$ouboul2012 . Once in orbit it was set at $\frac{9}{2}\,\nu_{orb}$ (V2 mode). In 2017 it was reported that at this spin rate the thermal noise of SUEP at the signal frequency of $0.92\,\rm{mHz}$ was about $2\times 10^{-10}\,\rm ms^{-2}/\sqrt{Hz}$ (Fig. 1, top right plot), while it was expected to be 10 times lower (CQG2019 , Fig. 11, top plot). In order to mitigate the problem the spin rate was increased to $\frac{35}{2}\,\nu_{orb}$ (V3 mode), moving the signal frequency from $0.92\,\rm mHz$ to $3.11\,\rm mHz$, with an expected noise reduction by a factor of $1.8$. Instead, the central plot of Fig. 1 shows that the reduction was by a factor of $4$, due to the fact that in going to spin mode V3 the best fitting line (2) jumped downwards (together with all fluctuations around it). At the new signal frequency $\nu_{{}_{EP}}=3.11\,\rm mHz$ the limiting noise is still a factor between 2 and 3 higher than expected (see CQG2019 , Fig. 11, top plot).
Figure 1: SUEP session #218, Slide 24 taken from the talk given by MICROSCOPE’s scientist M. Rodrigues in 2017 RodriguesBremen2017 . By increasing the spin frequency from mode V2 to mode V3 the frequency of the signal $\nu_{{}_{EP}}$ ($f_{{}_{EP}}$ in the slide) is moved to a higher value where noise is lower (as $1/\sqrt{\nu}$). This will ensure a noise reduction by a factor 1.8. However, it also happens that the best fitting line (2), along with fluctuations around it, undergoes a downward parallel translation to a lower level of noise at all frequencies, including the new $\nu_{{}_{EP}}=3.1\,\rm mHz$, such that the total noise reduction is by a factor of 4.
@@
It is apparent from (2) that the vertical position of the line (and of the plotted curve) is dictated by the acceleration $f_{TkQ}$, given by (3). It is hard to imagine that any of the parameters that $f_{TkQ}$ depends upon may have changed, in going from V2 to V3, by the amount required to explain the extra gain factor of about 2.2, never mentioned in later publications and unexplained to this date. Indeed, this event may not be exceptional, or due to the transition from lower to higher spin rate, since we observe very similar jumps even in sequential sessions at the same spin rate (see below).
@@
As the key limiting factor of each measurement, and of the final MICROSCOPE’s test of WEP, the quantity $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})=f_{TkQ}/\sqrt{2\pi\nu_{{}_{EP}}}$ deserves a careful scrutiny. For a given value of $f_{{}_{TkQ}}$ the measured noise is fitted by the same straight line. Once the oscillator has been set up and launched, $k$’s and $Q$’s are fixed, and (except for a mild dependence on $T$) the value of $f_{{}_{TkQ}}$ is fixed and so is the straight line (2) and its value at $\nu_{{}_{EP}}$.
@@
We cannot exclude a variation of the stiffness after a calibration, because in the calibration the quadratic acceleration term must be kept sufficiently small, which is achieved by offsetting the TM to a new zero, hence affecting the stiffness, $f_{{}_{TkQ}}$ and $S^{{}^{1/2}}(\nu_{{}_{EP}})$CatastiniCalibration . Conversely, in sequential sessions, or within the same session analyzed with equally valid methods, we expect no such variations. Should any anomalies occur, they will clearly appear in the values of $Q$ because $Q\propto 1/S(\nu_{{}_{EP}})$. Moreover, the extent and the relevance of the anomaly can be assessed by comparison with $Q$ measurements carried out in ground tests of WEP as well as in experiments for the measurement of the universal constant of gravity and the detection of gravitational waves.
@@
In order to carry out this analysis we need the values of $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ for all the valid sessions, but there is no table listing these values in the MICROSCOPE publications, as there is for $\delta_{i}\pm\sigma_{i}$, although accelerations are the measured physical quantities while $\delta$ is a derived number (see Eq. (6) TouboulCQG2022-9 ) and moreover $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ does not depend on the integration time –unlike $\delta$. With all $\delta$ values from Tables 6, 7 TouboulCQG2022-9 and the duration of each session from MissionScenarioCQG2022 we use (4) to derive all the corresponding values of $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$. The correctness of our derived values can be checked when the measured ones are published. [^4] A plot is shown in Fig. 11 ChhunCQG2022-5 of $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ for all sessions, including those which have not contributed to the final result of the WEP test, the purpose being to demonstrate an overall stability after the first few sessions of the mission. Although the numerical values of the plotted marks are not easy to read in the vertical scale used, for the 120-orbit SUEP session #236 (V3 mode, blue curve in Fig. 11, error bar shown on V2 mark by mistake) we can see that the plotted mark is much too high (by at least a factor of 3) with respect to the best fit value of about $2\times 10^{-11}\,\rm ms^{-2}/\sqrt{Hz}$ reported in Fig. 6 TouboulCQG2022-9 . By using (4) the plotted mark leads to a value of $\delta_{236}$ at least 3 times higher than reported in Table 7 TouboulCQG2022-9 , confirming the discrepancy. For other sessions of 120-orbit duration as #236, such as #238, #254 and #404 we know from (4) that the ratios of their $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ values to that of session #236 must be the same as the ratios of the corresponding values of $\delta$ (available in Table 7 TouboulCQG2022-9 ). While these ratios are never close to 1, their individual marks are all plotted at about the same height. in Fig. 11 ChhunCQG2022-5 . Thus, if this figure is correct, it contradicts Table 7 TouboulCQG2022-9 as well as Fig. 1 MS-PRL2022 and the final MICROSCOPE test of WEP published therein.
@@
Then, given the equilibrium temperature TouboulCQG2022-9 , the masses and the measured stiffnesses, the quality factor $Q$ is inferred from (3). Since $k_{1}\,,k_{2}$ may vary because of calibrations we compute also $k/Q$ assuming the same unknown $k$ for both TM’s. These quantities are listed in Table 1 for SUEP and in Table 2 for SUREF, along with the percentage $p$ of artificial data that have been introduced in each session after the elimination of glitches.
@@
Table 1: SUEP sensor unit. Session number in V2 or V3 spin mode; session $\delta$ and percentage of glitches $p$ from Tables 6,7 and 4,5 of TouboulCQG2022-9 ; calculated values of $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$, $Q$, $k/Q$. $\downarrow$ indicates a jump between sequential sessions; * an $M$ vs $A$ discrepancy (see text).
@@
For SUREF the 3 largest jumps are even larger, and in this case the sessions involved are all in V2 mode.
@@
Two methods of analysis, $M$-$ECM$ ($M$) and $ADAM$ ($A$), have been employed to estimate the experiment parameters, including $\delta$ and $\sigma$. Both methods use artificially reconstructed data to account for the missing data resulting from the elimination of glitches. [^5] Method $M$ estimates in the time domain the missing data, maximizing the likelihood conditional on the observed data, and then computes the least-squares estimate of the regression parameters. An estimation of the SD is also produced in the process. Method $A$ performs the parameter estimation in the frequency domain. As such it requires an uninterrupted, regularly spaced time series, which is obtained by filling the gaps left by the removal of glitches with the artificial data estimated by $M$. The two methods are stated to be equivalent, but in a few cases the values of $\delta$ calculated by $M$ and $A$ differ considerably (by a factor $\simeq 2$ in sessions #218, #438 and by a factor 7 in #442), implying that the same data stream leads to a different SD according to one or the other method. SUEP session #442 is an instructive extreme case. As reported in TouboulCQG2022-9 , Table 7, the noise is essentially the same but the values of $\delta$ differ by a factor 7; $\delta$ representing a level of WEP violation compatible with the measured acceleration difference noise. That is, for the same measurement session, in the same experimental conditions, the same data with the same level of noise lead to almost one order of magnitude difference in the evaluation of $\delta$. The difference is particularly evident when the results are interpreted in terms of the physics of the dominant thermal noise from internal damping (widely different values of $Q$, up to an unrealistic large value in session #442). SUREF session #778-1 is the only one, among all 32 sessions ($19$ SUEP $+$ $13$ SUREF), not affected by glitches, hence without artificial data. In this case, the results are exactly the same for $\delta$ and only slightly different for $\sigma$, as one expects if the same data set is analyzed with two equally valid methods (TouboulCQG2022-9 , Table 6). Sessions #442 and #778-1 show that the $M$ and $A$ methods are in agreement when applied to a time series of real measured data, but give different results (in particular different $\delta$ values) in the presence of artificial reconstructed data. Since the artificial data in $A$ are those estimated by $M$, it is the way the two methods manipulate these data that makes the difference.
Table 2: Same as Table 1 for SUREF sensor unit. A $\bullet$ has been added to indicate sessions with $Q$ values much larger than the largest value of $118$ measured in ground tests WT-RSI2000b .
@@
In SUEP session #218, the values of $\delta$ differ by a factor of about 2 depending on the type of analysis. This session (120 orbits) was the basis of the early results MS-PRL2017 , and was further elaborated two years later CQG2019 , ending up with a different value of the spectral density at the signal frequency. Even discarding the earlier results, it is apparent that the latest analysis cannot be considered conclusive. $Q$ values differing by a factor 3.5, for the same oscillator in the same conditions, are inexplicable.
The reference sensor SUREF behaves quite differently from SUEP. Thermal noise is about a factor $\sqrt{2}$ smaller than for SUEP (assuming the same $Q$), due to a factor $2$ smaller $(k_{1}/m_{1}^{2}+k_{2}/m_{2}^{2})$ term, and residual systematic errors are comparable to the stochastic errors TouboulCQG2022-9 . Four out of 9 sessions have been split because of sudden jumps in the mean value of the differential acceleration that required the two data segments to be treated as distinct experiments (in SUEP only 1 session, #326, out of 18 has been split). Unlike glitches, these jumps do not occur on all accelerometers simultaneously, hence they cannot be attributed to the spacecraft but are likely to originate in the accelerometers themselves, and last much longer (tens of seconds). In SUREF, in addition to the three cases of anomalous jumps mentioned above, we observe three sessions with large $M$ vs $A$ discrepancies (see Table 2). The values of $Q$ are typically larger than in SUEP and in three cases much larger than the largest value of $118$ reported in ground measurements WT-RSI2000b .
Early results MS-PRL2017 were based on session #176 (62 orbits). The spectral density $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ reported in MS-PRL2017 was confirmed in CQG2019 and attributed to $1/\sqrt{\nu}$ thermal noise, yielding $\delta=3.75\times 10^{-15}$ and $Q=65$, consistent with ground measurements. Instead, the final values of $\delta_{{M}}$ and $\delta_{{A}}$ are much smaller (Table 2), and the corresponding $Q$’s of 317 and 283 are far too high to be realistic on the basis of ground tests. Note that session #176 has $40\%$ of reconstructed data.
A question naturally arises: is there another source of noise with the same $1\sqrt{\nu}$ dependence like thermal noise from internal damping which may explain the anomalies described above?
As shown in Sec. II, the control of the TMs requires to control large acceleration biases. The acceleration bias $a_{bias}$ and the voltage $V_{bias}$ applied to control it are related through the scale factor $s_{{}_{f}}$:

### v4

@@ III Forces acting on Microscope accelerometers and their control
A higher stiffness results in a higher level of random thermal noise from internal damping. It also affects the electrostatic control of the TM which keeps it motionless at a selected capacitive zero so that the voltage applied yields (through a scale factor in $\rm ms^{-2}/V$ established by calibration) the acceleration which caused the motion (Sec. IV).
With drag free control on, the dominant forces acting on the test cylinders are due to the Earth tides and to the mechanical stiffness $k_{\parallel}$. No test cylinder is centered on the CoM of the spacecraft (Fig. 1), hence they are subjected to a tidal acceleration. If the orbit is circular and the attitude of the spacecraft is fixed in space, the tidal acceleration along the symmetry/sensitive $X$ axis reads:
$a_{tide}=\omega_{orb}^{2}\Big{[}X_{TM}\Big{(}\frac{3}{2}\cos{2\lambda}+\frac{1}{2}\Big{)}-\frac{3}{2}Z_{TM}\sin{2\lambda}\Big{]}$ (3)
where $\omega_{orb}=\sqrt{GM_{\oplus}/r^{3}}=2\pi\nu_{orb}$ is the orbital angular velocity, $\vec{\varrho}=(X_{TM},\,Z_{TM})$ is the instantaneous position vector of the TM relative to the CoM of the spacecraft in the orbit plane and $\lambda=\omega_{orb}t$ is the instantaneous longitude relative to the spacecraft-to-Earth direction. The tidal acceleration is $\propto 1/r^{3}$ and acts at frequency $2\nu_{orb}$ while a violation signal (at zero spin) is at $\nu_{orb}$. If the orbit has a non zero eccentricity $10<e<$ 1, the effect is larger at pericenter than at apocenter and there is also a tidal component at $\nu_{orb}$, like the signal, although reduced by the factor $e$. If the spacecraft is spinning, tidal effects are up-converted to $2\nu_{{}_{EP}}$ and $\nu_{{}_{EP}}$.
The difference of tidal accelerations between the two masses of each sensor unit is linear with the offsets $\Delta X,\,\Delta Z$ between their centers of mass within the SU. The largest effect at $2\nu_{EP}$ is then used to establish the values of these offsets, because tidal effects are deterministic and all the parameters they depend upon are measured, and this procedure allows a better determination than would be possible before launch. For SUEP the offsets measured in this way are $\Delta X\simeq 20\,\mu\rm m$ and $\Delta Z\simeq-5.6\,\mu\rm m$; for SUREF (in modulus) $\Delta Z$ is similar but $\Delta X$ is much larger ($\simeq-35.9\,\mu\rm m$).9 With SUEP measured values, the average tidal acceleration $\frac{1}{2}\omega_{orb}^{2}\Delta X$ as given in Eq.(3) is $\simeq 10^{-11}\,\rm ms^{-2}$.
While the tidal effect at $2\nu_{EP}$ can be easily separated from the signal, the one at $\nu_{EP}$ competes with it. It is reduced by the very small orbital eccentricity of Microscope ($e\simeq 10^{-4}$). In addition, as mentioned in Ref.(18), it is almost cancelled when the spacecraft is spinning. This can be achieved by an appropriate choice of the spin frequency (it is an odd multiple of half the orbital frequency) and of the initial phase of the sensitive axis relative to the Earth so that, at pericenter and apocenter, when the tidal effect is maximum or minimum –resulting in the variation at $\nu_{{}_{EP}}$– it is $90^{\circ}$ away from the direction to the Earth, and therefore it is not affected by the tide (Fig. 1, right sketch). This effect is indeed shown negligible.19
Under the tidal acceleration of Eq.(3) and the restoring elastic force due to $k_{\parallel}$ the TM will move, at any time, to an equilibrium position where the two effects balance each other, which occurs at an elongation $X_{eq}$ from the zero of the elastic force $0_{k}$. The electrostatic control acts to nullify this displacement and keep the TM at the selected capacitive zero $0_{C}$. If the two zeros do not coincide, the TM in $0_{C}$ is subjected to a constant elastic force that the control must nullify by applying a constant voltage (bias), in addition to the sinusoidal voltage required to nullify the tidal acceleration at $2\nu_{{}_{EP}}$.
@@
The resulting acceleration biases and their time evolution are reported, for all four accelerometers, in Ref.(15), Fig. 9. For SUEP the bias of the outer TM is $\simeq 10^{-6}\,\rm ms^{-2}$ and almost one order of magnitude smaller for the inner TM. With different offsets between their respective zeros, different masses and somewhat different stiffness, the acceleration bias is different for the two test masses of each sensor unit and the larger one dominates the acceleration difference. The time series for SUEP session #218 reported by Ref.(19), Slide 13, confirms that they are at the level of $10^{-6}\,\rm ms^{-2}$. For SUREF, the biases are of a few $10^{-7}\,\rm ms^{-2}$. 15
For such biases orders of magnitude above the target not to impair the measurement, it is mandatory that they remain constant. This will not be the case in the presence of patch charges on the surface of the TMs and their electrodes. Differences in the microscopic crystal structures, and impurities, over the surface of the test masses and their surrounding electrodes give rise to local potential variations and hence to spurious forces and torques (patch effects). The smaller the gap between the surfaces, the larger the effect. Random fluctuations of these potentials translate into a $1/\sqrt{\nu}$ spectral density. Moreover, occasional dislocations of these microstructures, whatever the cause, may produce sudden variations of the mean value of the fluctuations (Sec. IV).
Small force experiments require to take care of thermal effects. A well known “killer” in testing the WEP is the radiometer effect. Residual air pressure in the cage enclosing the TM, combined with a non zero temperature gradient along the symmetry axis, gives rise to a spurious acceleration along this axis which varies with the synodic frequency of the spacecraft relative to the Earth, it is inevitably different for the two masses, and therefore competes directly with a violation signal should the symmetry axis be –like in Microscope– the sensitive axis of the WEP test. The radiometer effect has been a matter of concern20 but direct in orbit measurements have shown that a very good thermal stability has been achieved (mostly by thermal insulation of the spacecraft), such that, at the $10^{-15}$ target level of Microscope, systematic thermal effects are smaller than random errors (SUEP) or close to them (SUREF).9; 8
IV Limiting noise, patch charges and unstable zero
@@
Ever since the preliminary Microscope results,5 the residual random acceleration differences in the frequency region of the signal were attributed to thermal noise from internal damping in the gold wires.
@@
In the case of internal damping the line which best fits the spectral density of the limiting noise decreases with the frequency as $1/\sqrt{\nu}$ and reads:21
@@
For a measurement session of duration $t_{int}$, as long as the acceleration noise is random, and therefore is bound to decrease as the square root of the integration time, the value of Eq.(4) at $\nu_{{}_{EP}}$ sets the limit $\delta$ to the violation signal established by the measurement session, according to the relationship:
@@
where $g_{{}_{drive}}=7.9\,\rm ms^{-2}$ is the gravitational acceleration from Earth at the spacecraft orbital distance, $\delta\cdot g_{drive}$ is a tiny acceleration difference between the TMs of the SU at the known signal frequency $\nu_{{}_{EP}}$, and $\delta$ is a solve-for parameter resulting from the confrontation between the predicted acceleration differences and their measured values (as defined in Ref.(8), Eq.(6)). [^1] Note that $\delta$ is never zero because thermal noise goes to zero only for an infinite measurement time. Should a WEP violation signal be buried by fluctuations around the best fitting line given by Eq.(4), the smaller their size, the shorter the additional integration time that would be required for it to emerge and be detected. These fluctuations yield the value $\sigma$ of the resulting $\delta\pm\sigma$ for the measurement session.
@@
Early on, the spin rate was designed to be between $3$ and $5$ times the orbital frequency $\nu_{orb}$.18 Once in orbit it was set at $\frac{9}{2}\,\nu_{orb}$ (V2 mode). In 2017 it was reported that at this spin rate the thermal noise of SUEP at the signal frequency of $0.92\,\rm{mHz}$ was about $2\times 10^{-10}\,\rm ms^{-2}/\sqrt{Hz}$ (Fig. 3, top right plot), while it was expected to be 10 times lower (Ref.(12), Fig. 11 top plot). In order to mitigate the problem the spin rate was increased to $\frac{35}{2}\,\nu_{orb}$ (V3 mode), moving the signal frequency from $0.92\,\rm mHz$ to $3.11\,\rm mHz$, with an expected noise reduction by a factor of $1.8$. Instead, the central plot of Fig. 3 shows that the reduction was by a factor of $4$, due to the fact that in going to spin mode V3 the best fitting line given by Eq.(4) jumped downwards (together with all fluctuations around it). At the new signal frequency of $3.11\,\rm mHz$ the limiting noise is still a factor between 2 and 3 higher than expected (Ref.(12), Fig. 11 top plot).
@@
It is apparent from Eq.(4) that the vertical position of the line (and of the plotted curve) is dictated by the acceleration $f_{TkQ}$, given by Eq.(5). The parameters that this acceleration depends upon can hardly have changed, in going from V2 to V3, by the amount required to explain the extra gain factor of about 2.2, never mentioned in later publications and unexplained to this date. Indeed, this event may not be exceptional, or due to the transition from lower to higher spin rate, since we observe very similar jumps even in sequential sessions at the same spin rate.
Figure 3: SUEP session #218, Slide 24 of Ref.(22). By increasing the spin frequency from mode V2 to mode V3 the frequency of the signal $\nu_{{}_{EP}}$ ($f_{{}_{EP}}$ in the slide) is moved to a higher value where noise is lower (as $1/\sqrt{\nu}$). This will ensure a noise reduction by a factor 1.8. However, it also happens that the best fitting line expressed by Eq.(4) undergoes a downward parallel translation (central plot) to a lower level of noise at all frequencies, including the new $\nu_{{}_{EP}}=3.11\,\rm mHz$, such that the total noise reduction is by a factor of 4 (unexplained).
@@
As the key limiting factor of each measurement, and of the final Microscope test of WEP, the quantity $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})=f_{TkQ}/\sqrt{2\pi\nu_{{}_{EP}}}$ deserves a careful scrutiny. For a given value of $f_{{}_{TkQ}}$ the measured noise is fitted by the same straight line. Once the oscillator has been set up and launched, $k$’s and $Q$’s are fixed, and (except for a mild dependence on $T$) the value of $f_{{}_{TkQ}}$ is fixed and so is the straight line of Eq.(4) and its value at $\nu_{{}_{EP}}$.
@@
We cannot exclude a variation of the stiffness after a calibration, because in the calibration the quadratic acceleration term must be kept sufficiently small, which is achieved by offsetting the TM to a new zero, hence affecting the stiffness, $f_{{}_{TkQ}}$ and $S^{{}^{1/2}}(\nu_{{}_{EP}})$.23 Conversely, in sequential sessions, or within the same session analyzed with equally valid methods, we expect no such variations. Should any anomalies occur, they will clearly appear in the values of $Q$ because $Q\propto 1/S(\nu_{{}_{EP}})$. Moreover, the extent and the relevance of the anomaly can be assessed by comparison with $Q$ measurements carried out in ground tests of WEP as well as in experiments for the measurement of the universal constant of gravity and the detection of gravitational waves.
@@
In order to carry out this analysis we need the values of $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ for all the valid sessions, but there is no table listing these values in the Microscope publications, as there is for $\delta_{i}\pm\sigma_{i}$,8 although accelerations are the measured physical quantities while $\delta$ is a derived number (see Ref.(8), Eq. (6)) and moreover $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ does not depend on the integration time –unlike $\delta$. With all $\delta$ values from Ref.(8), Tables 6, 7 and the duration of each session from Ref.(24) we use Eq.(6) to derive all the corresponding values of $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$. The correctness of our derived values can be checked when the measured ones are published. [^2] In Ref. (15), Fig. 11, the values of $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ are plotted for all sessions, including those which have not contributed to the final result of the WEP test, the purpose being to demonstrate an overall stability after the first few sessions of the mission. Although the numerical values of the plotted marks are not easy to read in the vertical scale used, for the 120-orbit SUEP session #236 (V3 mode, blue curve in Fig. 11, error bar shown on V2 mark by mistake) we can see that the plotted mark is much too high (by at least a factor of 3) with respect to the best fit value of about $2\times 10^{-11}\,\rm ms^{-2}/\sqrt{Hz}$ reported in Fig. 6, Ref.(8). The discrepancy is confirmed using Eq.(6), because the plotted mark leads to a value of $\delta_{236}$ at least 3 times higher than reported in Ref. (8), Table 7. For other sessions of 120-orbit duration like #236, such as #238, #254 and #404 we know from Eq.(6) that the ratios of their $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ values to that of session #236 must be the same as the ratios of the corresponding values of $\delta$ (available in Table 7 of Ref.(8)). While these ratios are never close to 1, their individual marks are all plotted at about the same height in Fig. 11 of Ref.(15). Thus, if this figure is correct, it contradicts Table 7 of Ref.(8) as well as Fig. 1 of Ref.(4) and the final Microscope test of WEP published therein.
@@
Then, given the equilibrium temperature,8 the masses and the measured stiffnesses, the quality factor $Q$ is inferred from Eq.(5). Since $k_{1}\,,k_{2}$ may vary because of calibrations we compute also $k/Q$ assuming the same unknown $k$ for both TM’s. These quantities are listed in Table 1 for SUEP and in Table 2 for SUREF, along with the percentage $p$ of artificial data that have been introduced in each session after the elimination of glitches.
@@
Table 1: SUEP sensor unit. Session number in V2 or V3 spin mode; session $\delta$ and percentage of glitches $p$ from Tables 6,7 and 4,5 of Ref.(8); calculated values of $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$, $Q$, $k/Q$. $\downarrow$ indicates a jump between sequential sessions; * an $M$ vs $A$ discrepancy (see text below).
@@
For SUREF the three largest jumps are even larger, and in this case the sessions involved are all in V2 mode.
@@
Two methods of analysis, $M$-$ECM$ ($M$) and $ADAM$ ($A$), have been employed to estimate the experiment parameters, including $\delta$ and $\sigma$. Both methods use artificially reconstructed data to account for the missing data resulting from the elimination of glitches. [^3] Method $M$ estimates in the time domain the missing data, maximizing the likelihood conditional on the observed data, and then computes the least-squares estimate of the regression parameters. An estimation of the SD is also produced in the process. Method $A$ performs the parameter estimation in the frequency domain. As such it requires an uninterrupted, regularly spaced time series, which is obtained by filling the gaps left by the removal of glitches with the artificial data estimated by $M$. The two methods are stated to be equivalent, but in a few cases the values of $\delta$ calculated by $M$ and $A$ differ considerably (by a factor $\simeq 2$ in sessions #218, #438 and by a factor 7 in #442), implying that the same data stream leads to a different SD according to one or the other method. SUEP session #442 is an instructive extreme case. As reported in Ref.(8), Table 7, the noise is essentially the same but the values of $\delta$ differ by a factor 7; $\delta$ representing a level of WEP violation compatible with the measured acceleration difference noise. That is, for the same measurement session, in the same experimental conditions, the same data with the same level of noise lead to almost one order of magnitude difference in the evaluation of $\delta$. The difference is particularly evident when the results are interpreted in terms of the physics of the dominant thermal noise from internal damping (widely different values of $Q$, up to an unrealistic large value in session #442). SUREF session #778-1 is the only one, among all 32 sessions ($19$ SUEP $+$ $13$ SUREF), not affected by glitches, hence without artificial data. In this case, the results are exactly the same for $\delta$ and only slightly different for $\sigma$, as one expects if the same data set is analyzed with two equally valid methods (Ref.(8), Table 6). Sessions #442 and #778-1 show that the $M$ and $A$ methods are in agreement when applied to a time series of real measured data, but give different results (in particular different $\delta$ values) in the presence of artificial reconstructed data. Since the artificial data in $A$ are those estimated by $M$, it is the way the two methods manipulate these data that makes the difference.
Table 2: Same as Table 1 for SUREF sensor unit. A $\bullet$ has been added to indicate sessions with $Q$ values much larger than the largest value of $118$ measured in ground tests. 17
@@
In SUEP session #218 the values of $\delta$ differ by a factor of about 2 depending on the type of analysis. This session (120 orbits) was the basis of the early results,5 and was further elaborated two years later,12 ending up with a different value of the spectral density at the signal frequency. Even discarding the earlier results, it is apparent that the latest analysis cannot be considered conclusive. $Q$ values differing by a factor 3.5, for the same oscillator in the same conditions, are inexplicable.
The reference sensor SUREF behaves quite differently from SUEP. Thermal noise is about a factor $\sqrt{2}$ smaller than for SUEP (assuming the same $Q$), due to a factor $2$ smaller $(k_{1}/m_{1}^{2}+k_{2}/m_{2}^{2})$ term, and residual systematic errors are comparable to the stochastic errors.8 Four out of nine sessions have been split because of sudden jumps in the mean value of the differential acceleration that required the two data segments to be treated as distinct experiments (in SUEP only one session, #326, out of 18 has been split). Unlike glitches, these jumps do not occur on all accelerometers simultaneously, hence they cannot be attributed to the spacecraft but are likely to originate in the accelerometers themselves, and last much longer (tens of seconds). In SUREF, in addition to the three cases of anomalous jumps mentioned above, we observe three sessions with large $M$ vs $A$ discrepancies (see Table 2). The values of $Q$ are typically larger than in SUEP and in three cases much larger than the largest value of $118$ reported in ground measurements.17
Early results 5 were based for SUREF on session #176 (62 orbits). The spectral density $S_{a}^{{}^{1/2}}(\nu_{{}_{EP}})$ reported in Ref.(5) was confirmed in Ref.(12) and attributed to $1/\sqrt{\nu}$ thermal noise, yielding $\delta=3.75\times 10^{-15}$ and $Q=65$, consistent with ground measurements. Instead, the final values of $\delta_{{M}}$ and $\delta_{{A}}$ are much smaller (Table 2), and the corresponding $Q$’s of 317 and 283 are far too high to be realistic on the basis of ground tests. Note that session #176 has $40\%$ of reconstructed data.
A question naturally arises: is there another source of noise with the same $1/\sqrt{\nu}$ dependence like thermal noise from internal damping which may explain the anomalies described above?
As shown in Sec. III, the control of the TMs requires to control large acceleration biases. The acceleration bias $a_{bias}$ and the voltage $V_{bias}$ applied to control it are related through the scale factor $s_{{}_{f}}$: